OpenStudy (anonymous):

If the lengths of two sides of a triangle are 5 and 9, respectively, which of the following could be the length of the third side of the triangle? a. 3 b. 5 c. 8 d. 15

OpenStudy (hybrik):

You can use the Pythagoras theorem, and else, where is the picture?

OpenStudy (anonymous):

no picture...

OpenStudy (hybrik):

Okay, then do 5^2 + 9^2 = c^2

OpenStudy (hybrik):

(AKA Pythagoras Theorem)

OpenStudy (anonymous):

doesn't work..

Nnesha (nnesha):

sum of two sides must greater than the 3rd one a+b >c b+c >a a+c> b

OpenStudy (carolinar7):

c^2=25+81 right?

OpenStudy (anonymous):

a+b<c<a-b

OpenStudy (campbell_st):

well there is a basic for triangles the sum of the shorter sides must be greater than the longest side... so if 9 is the longest and x is a short side you can say 5 + x > 9 if x is the longest side then 5 + 9 > x hope it helps

Nnesha (nnesha):

a = 1st side b= 3rd side c = 3rd sides

OpenStudy (hybrik):

You can do the 3rd side has to be bigger then the other sides

OpenStudy (anonymous):

got it thanks @Nnesha

Nnesha (nnesha):

yw

OpenStudy (anonymous):

So wait I thought I had it but I don't

OpenStudy (anonymous):

it can be all 4 options?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

only B and C

OpenStudy (anonymous):

Why?

Nnesha (nnesha):

first option a = 3 b = 5 c = 9 so a +b > c 3+ 5 > 9 is it correct?

OpenStudy (anonymous):

it is the reason that @campbell_st mentioned above

Nnesha (nnesha):

8 is greater than 9 ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

hey 15 works too right?

OpenStudy (anonymous):

but if you take 15, 5+9 < 15 so it doesnt work

OpenStudy (anonymous):

always any of the two sides should be equal or greater than the other

OpenStudy (anonymous):

ohhhhhhhh got it

Nnesha (nnesha):

:O what happened to ur post http://prntscr.com/7wm14b ?

Nnesha (nnesha):

http://i.imgur.com/ASDEMnt.png and i got mod power :o -.-

OpenStudy (anonymous):

lolol cool teehee

Nnesha (nnesha):

do you see this weird thing http://prntscr.com/7wm23v ?

Nnesha (nnesha):

nvm link isn't working

OpenStudy (anonymous):

yea on the first picture i do

Nnesha (nnesha):

ohhh weird ..... :/

OpenStudy (anonymous):

wow I didn't know you can have level 222 that's like equivalent to super saiyan level on here lmao

Nnesha (nnesha):

i was trying to delete ur post but eehh that was a dream

OpenStudy (anonymous):

lmaoo hey you can't delete me !!!! only i delete me!!!!!

Nnesha (nnesha):

lo alright gO_Od luck!

OpenStudy (anonymous):

i tried to close it but it's not working o.O

Nnesha (nnesha):

see ..

Nnesha (nnesha):

ohh closed congrats :D

OpenStudy (anonymous):

there u goooo I refreshed the page and it worked

Nnesha (nnesha):

thats what i did ..

OpenStudy (anonymous):

hmmm you know what?.... honestly i really love openstudy but the only thing i hate is that it lags wayyyyyy too much

Nnesha (nnesha):

ik but yeah it's oka nothing is perfect so ..

OpenStudy (anonymous):

yea